OpenStudy (vera_ewing):

How can (1/3)x – 2 = y and (1/4)x + 11 = y be set up as a system of equations? (5 points) 3y – x = –6 4y – x = 44 3y + x = –6 4y + x = 44 3y – 3x = –6 4y – 4x = 44 3y + 3x = –6 4y + 4x = 44

OpenStudy (vera_ewing):

@texaschic101 which one is it?

OpenStudy (vera_ewing):

A. 3y – x = –6 4y – x = 44 B. 3y + x = –6 4y + x = 44 C. 3y – 3x = –6 4y – 4x = 44 D. 3y + 3x = –6 4y + 4x = 44

OpenStudy (vera_ewing):

@jagr2713 please help!